# Thermal condensation with polybaric pressure

### v1

@@ 5.2.3 Velocity fluctuation and isobaric turbulence
In absence of any induced (external) large-scale velocity fluctuations, say from a supernova feedback term, the velocity field is developed self-consistently due to the formation of CNM and WNM. Besides, as our system is dissipation-less, we do not expect to see any shock front and the velocity is supposed to stay subsonic with respect to the local sound velocity at any point. In Fig.4, we show the local Mach number $M=v/\sqrt{T}$ (left panel, third row), the ratio of the net velocity to that of the local sound speed at a given point within the computational domain. As we can see, the local Mach number always stays well below $1.0$ in the entire computational domain and the plasma flow remains subsonic. We can compare this to the so-called turbulent Mach number $M_{t}$, determined through observation and defined as
@@
We note that the basic nature of thermal condensation mode is isobaric and the resultant nonlinear state should also show the nature of isobaric turbulence. We analyse the nature of these turbulence through a spatial power spectrum density (PSD) of the velocity field for both zero and strong magnetic field. In Fig.4, we show this spatial PSD for the case $\bm{B}>0,\beta=5$ (third row, right panel) and for $\bm{B}=0$ (bottom row, right panel) with respect to the spatial frequency (expressed as the inverse of the pixel of the velocity field image). What is shown is the 1D power spectrum density of the velocity field calculated using the Turbustat package (Koch et al., 2019). Note that for the 1D PSD, we have $P(k)\equiv E(k)\propto k^{-\kappa}$, where various values of the index $\kappa$ determine the nature of the turbulence. In the figure, the velocity PSDs are shown in the right panels of the third and fourth rows, respectively for the case when $\bm{B}>0$ (with $a_{p}=3$) and $\bm{B}=0$. In both these panels, the blue and red colored spectra (with relative error region marked) show the 1D PSD with a suggestive fitting (thick, solid, black line). The yellow colored spectra in both panels show the solenoidal part of the spectrum and the black colored part shows the compressional part (see the next paragraph for explanation). The left panel of the bottom row shows the compensated spectra for both cases (blue for $\bm{B}>0$ and red for $\bm{B}=0$) for more clarity.
@@
A few points should be noted while interpreting the power spectra shown in Fig.4. First, as our system is dissipation-less, the entire range of the power spectra is inertial as the so-called viscous dissipation range is practically zero. Which is why also, the spectra do not develop any marked transitions indicating the dissipation range. The spectra however show two distinct scalings, one which is marked with black solid line and the other with a green solid line in the figures. For the range shown in the figure by two vertical dashed lines, it is interesting to note that while the scalar pressure (zero $\bm{B}$) velocity field shows a scaling more akin to that of Burger with $E(k)\propto k^{-2}$, with a strong magnetic field $(\beta=5)$, the turbulence scales more like the classical incompressible Kolmogorov scaling with $E(k)\propto k^{-1.69}$. So, the effect of a magnetic field is to make the TI-induced isobaric turbulence more like an incompressible turbulence, at least at larger scales. It is also worthwhile to note that recent observations by Voyager I, which indicate incompressible turbulence in the very local interstellar medium (VLISM) (Zhao et al., 2020). To further confirm our findings, we have decomposed the PSD of the velocity field into two parts namely compressional and solenoidal parts through a Helmholtz-Hodge decomposition (Bhatia et al., 2013)
$\bm{v}=-\nabla\phi+\nabla\times A,$ (65)
where the $-\nabla\phi$ is the compressional part of the velocity field and the other solenoidal. From the figure, it is now amply clear that the total velocity PSDs closely follow their corresponding solenoidal parts for the most part which suggests that the flow is largely incompressible. However, there is marked difference between the two cases when $\bm{B}>0$ and $\bm{B}=0$. If we denote the deviation of the total PSD from the respective solenoidal PSD by $\Delta U_{k}$ for any $k$, we find that the mean deviation in case of scalar pressure $(\bm{B}=0)$ is twice as large as the same for the polybaric pressure $(\bm{B}>0)$ with $a_{p}=3$
$\left(\overline{\Delta U_{k}}\right)_{{\rm scalar}}\gtrsim 2\left(\overline{\Delta U_{k}}\right)_{{\rm polybaric}},$ (66)
which indicates that with magnetic field the flow becomes more incompressible, closely resembling Kolmogorov scaling. As mentioned, the turbulence remains largely isobaric in nature, which is confirmed by the pressure distribution function, shown in the bottom panel of Fig.5. The $\delta$ function-like behaviour of the pressure distribution function indicates that pressure remains isobaric. This turbulence, driven only by time-dependent heating and cooling, is surprisingly similar in nature to the turbulence driven by external forcing such as by supernova feedback or similar activities, creating a shock front, also known as the evaporation front (Kobayashi et al., 2022) in the WNM-CNM interface. Our analysis confirms this universal nature of turbulence in the ISM.
@@
which is $<l_{{\rm equiv}}$, the smallest structure that is generated in our simulation. So, we can see that viscosity plays only a minimal role in determining the overall structure formation due to TI.
It has been stated earlier that the presence of these two different scalings of the velocity PSD may be due to numerical diffusion (Kobayashi et al., 2022; Hennebelle & Audit, 2007). But, our analysis shows (as our code is specifically designed to take care of this very numerical diffusion, which is also the characteristics of our FCT code), we can safely argue that the presence of two scalings is not due to numerical diffusion, rather it indicates transition of the turbulence behaviour from either Kolmogorov’s or Burger’s at smaller scales.
@@ 5.2.4 Probability density function (PDF)
In Fig.5 (top row), we have shown the probability density function (PDF) for the scalar pressure and anisotropic pressure with $\beta=5$ and $10^{4}$ and $a_{p}=0$ and $3$ at $850\,{\rm Myr}$. As we can see, without any ambient magnetic field, density distribution shows clear indication of two distinctly separated phases with a modest amount of unstable gas in the intermediate region. However, with a background magnetic field, the growth of condensation is restricted, leading to a wide dispersion in the lower density regions. While for ‘weak’ magnetic field the condensations are roughly same for both $a_{p}=0$ and $3$, for the ‘strong’ magnetic field, we see a marked difference between the final condensation distribution; specifically, when $p_{\parallel}$ is considerably higher than $p_{\perp}$, the thermal condensation is slightly suppressed, resulting in a greater proportion of unstable gas compared to equal the case when $p_{\parallel}\sim p_{\perp}$. The middle panel of Fig.5 shows the PDF of thermal pressure (left), from which we can see that pressure-redistribution is quite affected by the magnetic field making the turbulence even more isobaric.
@@
 Refer to caption   Refer to caption
[…]
@@
In contrast to the above, the mass fraction in the WNM phase stays almost constant at about $\sim 16-17\%$ for both scalar and polybaric pressure, while that in the CNM phase decreases from about $\sim 67\%$ for scalar pressure to about $\sim 42\%$ for $a_{p}=3$ with $\beta=5$. The mass fraction in the UNM phase, however, increases by about $2.5$ times from $\sim 17\%$ for scalar pressure to about $\sim 41\%$ for $a_{p}=3$ with $\beta=5$. We note that all these findings are quite consistent with contemporary observational data, especially for the UNM phase (Heiles_2003; Kalberla & Haud, 2018; Murray et al., 2018). So, the presence of a background magnetic field considerably affects the mass and volume distribution of the WNM and CNM phases, with the magnetic field making the condensations more diffused. Many previous studies have neglected analysis of the long-lasting UNM phase, considering it to be a transient phase formed between the WNM and CNM phases. However, our analysis shows that UNM occupies a very significant amount of fractions both volume and mass-wise and turbulence remains crucial in maintaining this substantial quantity of thermally unstable gas. Similar natures of filling factors were also reported by earlier simulations (Kim et al., 2023; Kritsuk et al., 2017; Audit & Hennebelle, 2005).
[…]
@@
Our analysis of the three ISM phases further confirms the non-transient nature of the UNM and the volume and mass fractions of the multiphase ISM, calculated from our simulation is consistent to the contemporary results (Kim et al., 2023; Kritsuk et al., 2017; Audit & Hennebelle, 2005), which indicates to the fact that with a polybaric pressure profile, even without external forcing, we arrive at a pretty much same results, so far as the structure of the multiphase ISM is concerned. We also note that the background magnetic field makes the size of the WNM (in terms of volume) less than the case with scalar pressure, at the cost of the UNM, leaving the CNM size almost unchanged.
@@ 6 Summary and conclusion
This study examines the nonlinear development of isobaric TI due to time-dependent radiative heat-loss in a collisionless warm anisotropic and inviscid plasma using a 2D FCT simulation, for a polybaric equation of state across three different scenarios: unmagnetized, weakly magnetized, and strongly magnetized. The FCT simulation ensures that the simulation is not affected by numerical diffusion (Boris & Book, 1973). The simulation is started with a very small random perturbation in the density map, with an equilibrium, which is slightly unstable to the thermal condensation mode. As time progresses, the TI starts to build up with the linear phase, before entering a quasilinear and finally a nonlinear saturation regime. As such, we establish the linear instability conditions for various modes along with their corresponding growth rates (Field, 1965), which also serves as the benchmarking results for the nonlinear simulation. The effect of pressure anisotropy in the presence of an ambient magnetic field is handled with the polybaric equation of state (Stasiewicz, 2004), helping us to maintain the polytropic nature of the thermal pressure. We also show that for the parameter regime considered in this work, kinematic viscosity can be safely neglected. Apparently, our simulation with the relevant parameters, is able to resolve all the necessary time and length scales.
@@
  • The nature of the thermal instability-driven turbulence is similar in nature to the one driven by external forcing such as by supernova feedback and similar activities (Kobayashi et al., 2022; Kritsuk et al., 2017; Hennebelle & Audit, 2007; Saury et al., 2014). At the turbulent phase, most of the power in the PSD lies within the solenoidal mode, which confirms the incompressible nature of the flow. However, turbulence in the absence of a magnetic field surprisingly scales more like Burgers turbulence, even though no shock is involved in our simulation. In this case also, most of the power in PSD is solenoidal. Our analysis shows the universal nature of turbulent ISM, irrespective of any external forcing.
  • For the ‘strong’ magnetic field, the turbulence scales like the classical Kolmogorov turbulence, which indicates that turbulence remains largely incompressible. We also note similar kind of low frequency, isobaric, incompressible turbulence in the VLISM, as reported recently by Zhao et al. (2020).
@@
[…]
  • As far as the effect of polybaric pressure anisotropy is concerned, we have observed that the growth of TI is suppressed when the parallel component of the pressure becomes larger than the perpendicular component, leading to a comparatively larger fraction of the UNM. This is indicative of the fact that the anisotropy parameter $a_{p}$ plays a key role in the evolution of the ISM. Obviously, the effect of pressure anisotropy increases with the increase in magnetic field strength. For both ‘weak’ and ‘strong’ fields, the higher is the value of $a_{p}$, the lower is the mass and volume fraction of the condensation. However, the chosen strengths of the magnetic field do not show any pronounced difference in the evolution. We would also like to note that a stronger field is expected to cause defragmentation of the field aligned strands, which may be expected at a much lower $\beta$ for our case (Choi & Stone, 2012).
As a concluding remark, we would like to emphasize that, to our knowledge, this study is the first-ever comprehensive numerical study of TI in the ISM using a flux-corrected transport (FCT) method and also the first study to use polybaric pressure model in investigating TI in the ISM.
One of the authors, HS would like to thank CSIR-HRDG, New Delhi, India for the Senior Research Fellowship (File No: 09/059(0074)/2021-EMR-I)

### v2

@@ 5.3.3 Velocity fluctuation and isobaric turbulence
As our system is dissipation-less, we do not expect to see any shock front and the velocity is supposed to stay subsonic with respect to the local sound velocity at any point. This should not however be confused with supernovæ shocks, which we have assumed to have decayed into random fluctuations, considering the large-scale structure of the ISM that we are interested in. In Fig.7(e), we show the local Mach number $M=v/\sqrt{T}$, the ratio of the net velocity to that of the local sound speed at a given point within the computational domain. As we can see, the local Mach number always stays well below $1.0$ in the entire computational domain and the plasma flow remains subsonic. We can compare this to the so-called turbulent Mach number $M_{t}$, determined through observation and defined as
@@
We note that the basic nature of thermal condensation mode is isobaric and the resultant nonlinear state should also show the nature of isobaric turbulence. We analyse the nature of these turbulence through a spatial power spectrum density (PSD) of the velocity field for both zero and strong magnetic field. In Fig.7, we show this spatial PSD for the case $\bm{B}>0,\beta=5$ [Fig.7(f)] and for $\bm{B}=0$ [Fig.7(h)] with respect to the spatial frequency (expressed as the inverse of the pixel of the velocity field image). What is shown is the 1D power spectrum density of the velocity field calculated using the Turbustat package (Koch et al., 2019). Note that for the 1D PSD, we have $P(k)\equiv E(k)\propto k^{-\kappa}$, where various values of the index $\kappa$ determine the nature of the turbulence. In the figure, the velocity PSDs are shown in Figs.7(f, h), respectively for the case when $\bm{B}>0$ (with $a_{p}=3$) and $\bm{B}=0$. In both these panels, the blue and red colored spectra (with relative error region marked) show the 1D PSD with a suggested fitting (thick, solid, black line). The yellow colored spectra in both panels show the solenoidal part of the spectrum and the black colored part shows the compressional part (see the next paragraph for explanation). Fig.7(g) shows the compensated spectra for both cases (blue for $\bm{B}>0$ and red for $\bm{B}=0$) for more clarity.
@@
A few points should be noted while interpreting the power spectra shown in Fig.7. First, as our system is dissipation-less, the entire range of the power spectra in principle, is inertial as the so-called viscous dissipation range is practically zero. Which is why also, the spectra do not develop any marked transitions indicating the dissipation range. The spectra however show two distinct scalings with a knee, one which is marked with black solid line and the other with a green solid line in the figures. We believe that the existence of the latter one is due to the development of numerical viscosity in smaller scales. The so-called inertial range is shown in the figure by two vertical dashed lines. These PSD’s are with an initial velocity field having the ratio of power of the solenoidal part $(P_{\rm sol})$ to the compressive part $(P_{\rm com})$ as $(P_{\rm sol}:P_{\rm com})=(46:54)$, which is our Case (i). The final PSD for $\bm{B}>0$, however shows an almost equal solenoidal and compressive parts with a ratio of $(P_{\rm sol}:P_{\rm com})=(53:47)$. It is interesting to note that while the scalar pressure (zero $\bm{B}$) velocity field shows a scaling more akin to that of Burger with $E(k)\propto k^{-2}$, with a strong magnetic field $(\beta=5)$, the turbulence scales more close to the classical Kolmogorov scaling with $E(k)\propto k^{-1.69}$. Though the final stage has almost equal amounts of solenoidal and compressive powers in velocity, as expected from a weakly compressible component having low turbulent Mach number (see Fig.7), the solenoidal mode is expected to dominate (Elmegreen & Scalo, 2004), which is why we see an incompressible Kolmogorov-like scaling for the total PSD.
So, with a magnetic field, the TI-induced isobaric turbulence has grown to a nonlinear saturation with an almost equal amounts of solenoidal and compressive parts in velocity, with slightly more power in the solenoidal part, at least at larger scales. It is also worthwhile to note that recent observations by Voyager I, which indicate incompressible-like turbulence in the very local interstellar medium (VLISM) (Zhao et al., 2020). We also observe that in the presence of an ambient magnetic field, the final velocity PSD exhibits a Kolmogorov-like scaling, suggesting a transition toward more incompressible turbulence regime. As mentioned, the turbulence remains largely isobaric in nature, which is confirmed by the pressure distribution function, shown in the bottom panel of Fig.8. The $\delta$ function-like behaviour of the pressure distribution function indicates that pressure remains isobaric.
@@
which is $l_{{\rm equiv}}$, the smallest structure that is generated in our simulation. So, we can see that viscosity plays only a minimal role in determining the overall structure formation due to TI.
@@ 5.3.4 Probability density function (PDF)
In Figs.8(a) and (b), we have shown the probability density function (PDF) for the scalar pressure and anisotropic pressure with $\beta=5$ and $10^{4}$ and $a_{p}=0$ and $3$ at $850\,{\rm Myr}$. As we can see, without any ambient magnetic field, density distribution shows clear indication of two distinctly separated phases with a modest amount of unstable gas in the intermediate region. However, with a background magnetic field, the growth of condensation is restricted, leading to a wide dispersion in the lower density regions. While for ‘weak’ magnetic field the condensations are roughly same for both $a_{p}=0$ and $3$, for the ‘strong’ magnetic field, we see a marked difference between the final condensation distribution; specifically, when $p_{\parallel}$ is considerably higher than $p_{\perp}$, the thermal condensation is slightly suppressed, resulting in a greater proportion of unstable gas compared to the case when $p_{\parallel}\sim p_{\perp}$. Whereas the WNM mass fraction remain almost same $(16\%-17\%)$ for $a_{p}=0$ to $3$, CNM mass fraction reduces from $49\%$ to $42\%$ and in turn UNM mass fraction increases from $34\%$ to $41\%$. Detailed procedure of these analyses is mentioned in the following paragraph. Fig.8(c) shows the PDF of thermal pressure, from which we can see that pressure-redistribution is quite affected by the magnetic field making the turbulence even more isobaric.
@@
Figure 8: Mass-weighted density PDFs for $\beta=10^{4}$ [panel (a)] and $\beta=5$ [panel (b)]. Panel (a) also shows the density PDF for scalar pressure. Panel (c) shows the PDFs for pressure and panel (d) shows the PDFs for temperature. The bottom row shows the volume fraction $(\%)$ of the three phases across evolution timeline for scalar pressure [panel (e)] and for $\beta=5,a_{p}=3$ [panel (f)].
Figure 9: Left: A velocity field with a dominant solenoidal component $(\sim 72\%)$ and a minor compressive component $(\sim 28\%)$, and Right: A predominantly solenoidal component $(\gtrsim 95\%)$ and a negligible compressive component.
We have also calculated the volume fraction occupied in the temperature PDFs, considering the unstable log–temperature range of 0.7 0.25-0.7\,-\,0.25, which corresponds to physical temperature $\sim 100-890\,{\rm K}$. We find that minimum and maximum temperatures indicated in our simulation results lie between $\sim 65-6400\,{\rm K}$, which are within the limits of observed temperatures for WNM, UNM, and CNM (Heiles & Troland, 2003; Stahler & Palla, 2004). In the bottom row of Fig.8, we have shown the volume fraction $(\%)$ occupied over time for the three phases – WNM, CNM, and UNM for scalar pressure [Fig.8(e)] and for $\beta=5\,\textrm{with}\,a_{p}=3$ [Fig.8(f)]. The dashed vertical lines in both panels indicate a transition of the TI from linear to nonlinear regime. We have calculated the volume and mass fractions occupied by these three phases at $t=850\,{\rm Myr}$ from the density PDFs. We consider the log–density range $\sim-0.25-0.7$ to be the unstable phase range which corresponds to density $\sim 1-10\,{\rm cm^{-3}}$. Densities below this range are considered to be WNM and above are considered to be CNM. We find that while the volume fraction in WNM decreases from about $\sim 80\%$ for scalar pressure to about $\sim 68\%$ for $a_{p}=3$ with $\beta=5$, the same for CNM stays almost constant at about $\sim 9-7\%$. Naturally, we have a considerable portion of the volume fraction in the UNM phase which increases from about $\sim 11\%$ for scalar pressure to almost two times of that $\sim 25\%$ for $a_{p}=3$ with $\beta=5$.
@@
In contrast to the above, the mass fraction in the WNM phase stays almost constant at about $\sim 16-17\%$ for both scalar and polybaric pressure, while that in the CNM phase decreases from about $\sim 67\%$ for scalar pressure to about $\sim 42\%$ for $a_{p}=3$ with $\beta=5$. The mass fraction in the UNM phase, however, increases by about $2.5$ times from $\sim 17\%$ for scalar pressure to about $\sim 41\%$ for $a_{p}=3$ with $\beta=5$. We note that all these findings are quite consistent with contemporary observational data, especially for the UNM phase (Heiles & Troland, 2003; Kalberla & Haud, 2018; Murray et al., 2018). So, the presence of a background magnetic field considerably affects the mass and volume distribution of the WNM and CNM phases, with the magnetic field making the condensations more diffused. Many previous studies have neglected analysis of the long-lasting UNM phase, considering it to be a transient phase formed between the WNM and CNM phases. However, our analysis shows that UNM occupies a very significant amount of fractions both volume and mass-wise and turbulence remains crucial in maintaining this substantial quantity of thermally unstable gas. Similar natures of filling factors were also reported by earlier simulations (Kim et al., 2023; Kritsuk et al., 2017; Audit & Hennebelle, 2005).
Regarding the distribution of size, we find that at $500\,{\rm Myr}$, the average size of cloud condensates without a magnetic field is $\sim(6\times 2)\,{\rm pc^{2}}$ and for a background magnetic field, the average filament size is $\sim(14\times 0.16)\,\text{pc}^{2}$ along the field direction. It is to be noted that these values are consistent with observations of ISM HI. For example, McClure-Griffiths et al. (2006) observed a network of dozens of hair-like filaments of magnetically dominated cold hydrogen with lengths up to $\sim 17\,{\rm pc}$ and widths $0.1\,{\rm pc}$ while Clark et al. (2014) have observed magnetically aligned HI fibres of length $\sim 8.7\,\text{pc}$ and width $0.12\,\text{pc}$ by studying diffuse galactic interstellar medium.
@@
Our analysis of the three ISM phases further confirms the non-transient nature of the UNM and the volume and mass fractions of the multiphase ISM, calculated from our simulation is consistent to the contemporary results (Kim et al., 2023; Kritsuk et al., 2017; Audit & Hennebelle, 2005), which indicates to the fact that with a polybaric pressure profile, we arrive at a pretty much same results, so far as the structure of the multiphase ISM is concerned. We also note that the background magnetic field makes the size of the WNM (in terms of volume) less than the case with scalar pressure, at the cost of the UNM, leaving the CNM size almost unchanged.
Table 1: Comparison of final volume and mass fractions
Volume fractions (%)
Initial velocity field UNM CNM WNM
(i)   $P_{{\rm com}}\sim 54\%,P_{\rm sol}\sim 46\%$   $29$ 66 $65$
(ii)  $P_{{\rm sol}}\sim 72\%,P_{{\rm com}}\sim 28\%$   $35$ 55 $60$
(iii) $P_{{\rm sol}}\gtrsim 95\%$   $30$ 66 $64$
Mass fractions (%)
(i)   $P_{{\rm com}}\sim 54\%,P_{\rm sol}\sim 46\%$   $47$ $36$ $17$
(ii)  $P_{{\rm sol}}\sim 72\%,P_{{\rm com}}\sim 28\%$   $55$ $26$ 1919
(iii) $P_{{\rm sol}}\gtrsim 95\%$   $48$ $33$ 1919
5.3.5 Effect of a solenoidal velocity field
Finally, in this section, we present the comparative results of an initial velocity field with a dominant solenoidal component. We consider the other two cases, as mentioned in Section 5.1. In Fig.9, we have shown the PSDs for these two cases – Case (ii): one with the ratio of $\sim 72:28$ for the solenoidal and compressive parts and Case (iii): one with a dominant solenoidal part with only a negligible compressive component.
The volume and mass fractions of the resultant multiphase ISM for an initial velocity field with dominant solenoidal component is shown in Fig.10. In the figure, the left and right panels, respectively correspond to the Cases (ii) and (iii) discussed in Section 5.1. The top row [Figs.10(a, b)] shows the volume fractions and the bottom row [Figs.10(c, d)] shows the mass fractions. These volume fractions should be compared to the one with a velocity field with a relatively dominant compressive field shown in Fig.8. Apart from a relatively earlier starting time for growth of the nonlinear phase and subsequent saturation, statistically the volume and mass fractions of the WNM, UNM, and CNM for these two cases are similar to one with a relatively dominant compressive velocity field. The comparative values of the volume and mass fractions of these cases are shown in Table. 1, where the quantity $P_{\rm com,sol}$ denotes, respectively, the power in compressive and solenoidal parts of the initial velocity field. In all these cases, the volume and mass fractions are estimated during the quasi-steady state of the evolution when the nonlinear TI has saturated, which is about $\sim 1000\,\rm Myr$ for Case (i) and $\sim 1150\,\rm Myr$ for Cases (ii) and (iii). From the table, we can infer that a more solenoidal initial velocity field (Case (ii)) favours a larger fraction of UNM at the cost of CNM and WNM, both mass and volume wise. Further increasing the solenoidal component in Case (iii) reduces the mass and volume fractions, bringing them closer to their original values in Case (i).
Regarding the timeline of development of the TI, we see that while for a relatively dominant compressive velocity field (with a marginally weak solenoidal part), the separation of the multiphase ISM starts around $450\,\rm Myr$, this begins at around $650-700\,\rm Myr$ for a velocity field with dominant solenoidal part. We note that in all these cases, the average amplitude of the initial velocity field is same, so the late begin of the separation of ISM phases in the latter can be attributed to the fact that a compressive perturbation helps drive the TI making it reach its nonlinear phase in a comparatively short time. We must however emphasize that the absolute time when the nonlinear phase begins which makes the beginning of the multiphase separation process is dependent on the initial perturbation strength – the higher the initial perturbation strength, the quicker the TI reach its nonlinear saturation. For the purpose of comparison, the mass fractions for the multiphase ISM for a relatively dominant compressive initial velocity field, a case which we have already discussed in the previous sections, is shown in Fig.11.
Analysis of the spatial PSD at the final stage of thermal instability reveals that the inertial-range turbulence scaling steepens from $E(k)\propto k^{-1.35}$ to $E(k)\propto k^{-1.43}$ as the initial velocity field changes from moderately solenoidal to predominantly solenoidal, respectively. The effects of this are evident in the multiphase segregation and mass and volume fractions discussed earlier. The steeper scaling observed in Case (iii) suggests a more efficient energy cascade, resulting in enhanced phase segregation compared to Case (ii), and finally leading to a higher CNM fraction.
Table 2: Ratio of power $(P)$ in the initial and final velocity fields for different cases.
Case Initial power ratio Final power ratio
$(P_{{\rm sol}}:P_{{\rm com}})$ $(P_{{\rm sol}}:P_{{\rm com}})$
(i) $46:54$ $53:47$
(ii) $72:28$ $52:48$
(iii) $95:5$ $55:45$
We also note the powers of velocity fields in the saturation stage of the instability, separate to an almost equal amounts in the solenoidal and compressive parts with a slightly more power in the solenoidal part, irrespective of the initial velocity field (see Table.2).
Figure 10: Volume (top row) and mass (bottom row) fractions of multiphase ISM for an initial velocity field with a dominant solenoidal component $(\sim 72\%)$ and a minor component $(\sim 28\%)$ [panels (a), (c)], and a predominantly solenoidal component $(\gtrsim 95\%)$ and a negligible compressive component [panels (b), (d)]. These two cases correspond to the corresponding velocity fields shown in Fig.9.
@@ 6 Summary and conclusion
This study examines the nonlinear development of isobaric TI due to time-dependent radiative heat-loss in a collisionless warm anisotropic and inviscid plasma using a 2D FCT simulation, for a polybaric equation of state across three different scenarios: unmagnetized, weakly magnetized, and strongly magnetized. The FCT simulation ensures that numerical diffusion remains minimised throughout the simulation while preventing unphysical oscillations (Boris & Book, 1973). The simulation is started with a random velocity field with an adjustable compressive and solenoidal parts, which is supposed to be a result of supernovæ forcing. As time progresses, the TI starts to build up with the linear phase, before entering a quasilinear and finally a nonlinear saturation regime. As such, we establish the linear instability conditions for various modes along with their corresponding growth rates (Field, 1965), which also serves as the benchmarking results for the nonlinear simulation. The effect of pressure anisotropy in the presence of an ambient magnetic field is handled with the polybaric equation of state (Stasiewicz, 2004), helping us to maintain the polytropic nature of the thermal pressure. We also show that for the parameter regime considered in this work, kinematic viscosity can be safely neglected. Apparently, our simulation with the relevant parameters, is able to resolve all the necessary time and length scales.
@@
  • In the absence of background magnetic field, a relatively dominant compressive forcing drives the turbulence to Burger like inertial range scaling but with relatively more power in solenoidal component of the PSD. The nature of the turbulence remains subsonic.
  • For the ‘strong’ magnetic field with relatively dominant compressive forcing, the turbulence scales like the classical Kolmogorov turbulence, suggesting a transition toward an incompressible, subsonic turbulent regime. We also note similar kind of low frequency, isobaric, incompressible turbulence in the VLISM, as reported recently by Zhao et al. (2020). We have also observed that, for moderately dominant solenoidal forcing, phase segregation relatively gets suppressed a little due to inefficient turbulence cascading as confirmed by a flatter inertial range scaling in PSD. In contrast to this, under predominantly solenoidal forcing, the relatively steeper inertial-range scaling, approaching the Kolmogorov 5/3-5/3 slope, suggests a more efficient energy cascade leading to a higher fraction of CNM in the final stage.
@@
    Figure 11: The mass fractions of multiphase ISM for a relatively dominant compressive initial velocity field. This for the case that we have already discussed in Section 5.2. This case is for $\beta=5$ and polybaric parameter $a_{p}=3$. The corresponding volume fraction is shown in Fig.8(f).
  • The presence of a homogenous background magnetic field severely affects the formation of condensations, mainly restricting the fluid flow in the perpendicular direction and thereby elongating the condensates to form filaments along the direction of the magnetic field with an average length of about $\sim 14\,\text{pc}$ and width $\sim 0.16\,\text{pc}$, which closely agrees with existing observational studies (McClure-Griffiths et al., 2006; Clark et al., 2014).
  • As far as the effect of polybaric pressure anisotropy is concerned, we have observed that the growth of TI is suppressed when the parallel component of the pressure becomes larger than the perpendicular component, leading to a comparatively larger fraction of the UNM. This is indicative of the fact that the anisotropy parameter $a_{p}$ plays a key role in the evolution of the ISM. Obviously, the effect of pressure anisotropy increases with the increase in magnetic field strength. For both ‘weak’ and ‘strong’ fields, the higher is the value of $a_{p}$, the lower is the mass and volume fraction of the condensation. However, the selected magnetic field strengths do not exhibit any significant differences in the overall evolution, apart from a small enhancement of anisotropic effects. We would also like to note that a stronger field is expected to cause defragmentation of the field aligned strands, which may be expected at a much lower $\beta$ for our case (Choi & Stone, 2012).
  • The formation of condensations is strongly influenced by both thermal conduction and pressure anisotropy. Without thermal conduction, filamentary structures tend to form across the direction of magnetic field (Wareing et al., 2016), whereas the presence of conduction aligns the condensations along the magnetic field direction. Additionally, stronger magnetic fields enhance the effects of pressure anisotropy. Higher anisotropy leads to filaments getting elongated along the direction transverse to the magnetic field, while lower anisotropy results in more scattered, nearly isotropic structures.
As a concluding remark, we would like to emphasize that, to our knowledge, this study is the first-ever comprehensive numerical study to use polybaric pressure model in investigating TI in the ISM, thereby offering new insights regarding the effects of pressure anisotropy on multiphase structure formation and we believe it establishes a foundation for future investigations, which should address the case of tangled magnetic fields evolution for a more realistic system.
Acknowledgements
One of the authors, HS would like to thank CSIR-HRDG, New Delhi, India for the Senior Research Fellowship grant no: 09/059(0074)/2021-EMR-I. The authors to like to thank the referee for constructive suggestions.
Data Availability
The data that support the findings of this study are available within the article.